How do you graph the circle x squared + y squared + 4x + 6y - 3 = 0?
Answer by stanbon(75887)

How do you graph the circle x squared + y squared + 4x + 6y - 3 = 0?
------
Complete the square on the x-terms and on the y-terms:
x^2 + 4x + 4 + y^2 + 6y +9 = 3+4+9
------
Factor:
(x+2)^2 + (y+3)^2 = 16
----
Sketch a circle with:
center at (-2,-3)
radius = sqrt(16) = 4
==========================
Cheers,
Stan H.